# Inference on many jumps in nonparametric panel regression models

### v2

@@ 3.1 Assumptions
Due to the construction of $\varepsilon_{j}(\cdot,\cdot)$ in (2.1), if functions $h_{j}(\cdot,\cdot)$ and $\tau_{j}(\cdot,\cdot)$ are both bounded, then the above assumption naturally holds. Let $\bar{b}=\max_{1\leq j\leq N}b_{j}$ and $\underline{b}=\min_{1\leq j\leq N}b_{j}.$ The next assumption imposes some smoothness conditions.
Assumption 3.4.
@@
Assumption 3.4 $(i)$ imposes a lower bound condition for the density functions $g_{j}(\cdot)$. Condition $(ii)$ imposes some smoothness conditions of the mean function $f_{j}(\cdot).$ Note that we do not require $f_{j}(\cdot)$ to be differentiable at point $x$; it suffices for $f_{j}(\cdot)$ to have bounded left and right second derivatives near $x$. Assumption $(iii)$ is for the smoothness of the variance function $\sigma_{j}^{2}(\cdot,\cdot)$. If $\sigma_{j}(x,u)$ has bounded partial left and right derivatives with respect to $x$, for any $x\in\mathcal{C}_{j}$ uniformly over $u,$ then assumption $(iii)$ is fulfilled. Note that we do not assume the left and right derivatives of $f_{j}(\cdot)$ to be the same, meaning the slope around the threshold can differ. To be general, we do not impose any differentiability assumptions on $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$.            $\Box$
@@
Assumption 3.5$(i)$-$(ii)$ imposes certain conditions on the kernel function $K\left(\cdot\right)$and bandwidth parameters $b_{j}$’s, respectively. If one chooses $b_{j}$ to be the same across $j,$ Assumption 3.5$(ii)$ reduces to the standard conditions on the bandwidth parameter in local linear kernel estimation with one-dimensional covariate.            $\Box$
Let $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1})$ be the conditional joint density of $X_{j_{1}t}$ and $X_{j_{2}t}$ given $\mathcal{F}_{t-1}.$ The following assumption imposes a condition on $g_{j_{1},j_{2}}(\cdot,\cdot|\mathcal{F}_{t-1}).$
Assumption 3.6.
3.2 Gaussian Approximation Results
If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in subsections 3.2 and 3.3. We first consider the GA result for $\mathcal{I}$ defined in (2.13). Define $d_{j}=\gamma_{j}/v_{j}$ and $\underline{d}=(d_{1},d_{2},\ldots,d_{N})^{\top}$. Further define $Z+\underline{d}$ as a Gaussian vector centered around $\underline{d}$, where $Z$ is a vector of $N$ standard normally distributed random variables, and $\underline{d}$ denotes a bias term which is a zero vector under the null $H_{0}^{\left(1\right)}$. The following theorem states that the infeasible statistic $\mathcal{I}$ can be approximated well by the maximum of Gaussian random variables. Denote
Theorem 3.1 (Gaussian Approximation 1).
Let Assumptions 3.1-3.6 hold. Then we have
Now we list the related rate requirement under different moment conditions. Under Assumption 3.1$(i)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $\underline{b}^{-1}T^{2/q-1}\mathrm{log}^{3}(NT)\rightarrow 0$ and $\mathrm{log}(NT)N^{1/q}T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$. Under Assumption 3.1$(ii)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $(T\underline{b})^{-1}\mathrm{log}^{7}(NT)\rightarrow 0,$ $\bar{b}\mathrm{log}^{3}(N)\rightarrow 0$ and $\mathrm{log}(NT)T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$.
Theorem 3.1 lays down the foundation for testing the null hypothesis $H_{0}^{\left(1\right)}.$ The proof of Theorem 3.1 is technically challenging and involved for two reasons. First, we allow for both temporal and cross-sectional dependence among $\left\{\left(X_{jt},\epsilon_{jt}\right)\right\}.$ Second, the latent variable $U_{jt}$ makes the error term $e_{jt}$ in (1.1) highly complex. Such complicated structure makes it impossible to apply any existing GA results or their proof strategies directly to prove the above result. In Section A.1 of the online supplement, we outline the main idea used in the proof of the above theorem, highlighting that the key step is proving the GA result in Lemma A.1. We note that the term $\mathrm{log}^{7/6}(NT)(\underline{b}T)^{-1/6}+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}$ arises from extending the original GA results in Chernozhukov et al., 2017a . The term $\mathrm{log}(NT)N^{1/q}T^{-\beta}$ is due to the temporal dependence and the moment condition of the innovation term $\eta_{jt}$. The term $\bar{b}^{1/3}\mathrm{log}^{2/3}(N)$ results from the covariance approximation and Gaussian comparison.
$\approx\max_{1\leq j\leq N}(Tb_{j})^{1/2}\Big{|}\sum_{t=1}^{T}\Big{(}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}-\sum_{j=1}^{N}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}/N\Big{)}/\tilde{v}_{j}+\tilde{d}_{j}\Big{|},$ (3.6)
where $\tilde{d}_{j}=(\gamma_{j}-\frac{1}{N}\sum_{i=1}^{N}\gamma_{i})/\tilde{v}_{j}.$ The approximation is valid due to the smoothness of $f_{j}(\cdot)$ and the remainder is addressed in Subsection B.3. Let $\tilde{\underline{d}}=(\tilde{d}_{1},\tilde{d}_{2},\ldots,\tilde{d}_{N})^{\top}.$ The following theorem states that the infeasible statistic $\mathcal{Q}$ can be approximated well by the maximum of Gaussian random variables.
Theorem 3.2 (Gaussian Approximation 2).
Suppose that Assumptions 3.1-3.6 hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\Big{(}\mathcal{Q}\leq u\Big{)}-\mathbb{P}(|Z+\tilde{\underline{d}}|_{\infty}\leq u)\Big{|}\lesssim\mathcal{R}_{NT}.$ (3.7)
Theorem 3.2 establishes the foundation for testing the null hypothesis $H_{0}^{\left(2\right)}.$ Its proof cannot directly follow from Theorem 3.1 due to the introduction of the additional averaging term of $Y_{jt}$ which might not be of smaller order. However, the steps can be simplified by imposing weak cross-sectional dependence, as the additional averaging term would then be of higher order under this assumption.

### v3

@@ 3.1 Assumptions
Let
$\bar{b}=\max_{1\leq j\leq N}b_{j}\quad\text{and}\quad\underline{b}=\min_{1\leq j\leq N}b_{j}.$ (3.4)
Assumptions SA-1-SA-3 in Section A.1 in the Supplementary Material are concerning standard assumptions on boundedness, smoothness and kernel functions involved.
3.2 Gaussian Approximation Results
If the true threshold is given, then $\mathcal{C}_{j}=\{c_{0j}\}$ is a single point in Subsections 3.2 and A.3. We first consider the GA result for $\mathcal{I}$ defined in (2.13). Define $d_{j}=(Tb_{j})^{1/2}\gamma_{j}/v_{j}$ and $\underline{d}=(d_{1},d_{2},\ldots,d_{N})^{\top}$. Further define $Z$ as a centered Gaussian random vector with identity covariance matrix. Under the null $H_{0}^{\left(1\right)},$ the bias term $\underline{d}$ becomes a zero vector. The following theorem states that the infeasible statistic $\mathcal{I}$ can be approximated well by the maximum of Gaussian random variables. Denote
Theorem 3.1 (Gaussian Approximation 1).
Let Assumptions 3.1-3.3 and SA-1-SA-3 hold. Then we have
Now we list the related rate requirement on $T$, $\bar{b}$ and $N$ under different moment conditions. Under Assumption 3.1$(i)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $\underline{b}^{-1}T^{2/q-1}\mathrm{log}^{3}(NT)\rightarrow 0$ and $\mathrm{log}(NT)N^{1/q}T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$. Under Assumption 3.1$(ii)$, if $T\bar{b}^{5}\mathrm{log}(N)\rightarrow 0,$ $(T\underline{b})^{-1}\mathrm{log}^{7}(NT)\rightarrow 0,$ $\bar{b}\mathrm{log}^{3}(N)\rightarrow 0$ and $\mathrm{log}(NT)T^{-\beta}\rightarrow 0,$ then $\Delta\rightarrow 0$ and $\mathcal{R}_{NT}\to 0$.
Theorem 3.1 lays down the foundation for testing the null hypothesis $H_{0}^{\left(1\right)}.$ The proof of Theorem 3.1 is technically challenging and involved for two reasons. First, we allow for both temporal and cross-sectional dependence among $\left\{\left(X_{jt},\epsilon_{jt}\right)\right\}.$ Second, the latent variable $U_{jt}$ makes the error term $e_{jt}$ in (2.1) highly complex. Such complicated structure makes it impossible to apply any existing GA results or their proof strategies directly. In Section B.1.1 of the online supplement, we outline the main idea used in the proof of the above theorem, highlighting that the key step is proving the GA result in Lemma B.1. We note that the term $\mathrm{log}^{7/6}(NT)(\underline{b}T)^{-1/6}+\mathrm{log}^{1/2}(N)T^{1/2}\bar{b}^{5/2}$ arises from extending the original GA results in Chernozhukov et al., 2017a . The term $\mathrm{log}(NT)N^{1/q}T^{-\beta}$ is due to the temporal dependence and the moment condition of the innovation term $\eta_{jt}$. The term $\bar{b}^{1/3}\mathrm{log}^{2/3}(N)$ results from the covariance approximation and Gaussian comparison. Here we adopt maximum based test statistic, extending to other type of statistic can also be considered, see for example the statistics in Fan et al., (2015) and Giessing and Fan, (2020).
$\approx\max_{1\leq j\leq N}\Big{|}(Tb_{j})^{1/2}\sum_{t=1}^{T}\Big{(}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}-\sum_{j=1}^{N}(w_{jt,b}^{+}-w_{jt,b}^{-})e_{jt}/N\Big{)}/\tilde{v}_{j}+\tilde{d}_{j}\Big{|},$ (3.6)
where $\tilde{d}_{j}=(Tb_{j})^{1/2}(\gamma_{j}-\frac{1}{N}\sum_{i=1}^{N}\gamma_{i})/\tilde{v}_{j}.$ The approximation is valid due to the smoothness of $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$, the remainder is addressed in Subsection C.3. Let $\tilde{\underline{d}}=(\tilde{d}_{1},\tilde{d}_{2},\ldots,\tilde{d}_{N})^{\top}.$ We deliver the GA results for $\mathcal{Q}$ in Supplementary Material A.2 and the proofs in Section B.2. Furthermore, the theoretical results regarding the derivative are given in Section A.5 and the proofs in Section B.7. In addition, in Section A.3, we establish the validity of the test statistics when variance estimators are utilized.
The most interesting finding is that in the GA result (3.5), neither temporal nor cross-sectional dependencies appear in $|Z+\underline{d}|_{\infty}$, despite allowing strong cross-sectional dependence for the underlying data generating process $X_{jt},U_{jt},\epsilon_{jt}$ within $\mathcal{I}$. The theoretical intuition behind this result is that local smoothing effectively whitens the noise. In practice, this observation suggests that, when the Assumption 3.3 hold for cross-sectional dependence, the correlation in the estimated statistics can be safely ignored, requiring only adjustments for heteroskedasticity.
3.3 Theoretical Results for the Unknown Threshold Case
In this section, we present theoretical results in the case of unknown thresholds. For simplicity of notation, we define $\mathcal{C}_{j}$ as the unified interval $\mathcal{C}_{j}=[c_{\min},c_{\max}]$, where $c_{\min}$ and $c_{\max}$ are constants in $\mathbb{R}$. The theorem still holds even when $\mathcal{C}_{j}$ differs for different $j$. Recall that $[c_{1},\ldots,c_{K}]$ is the grid we search over. Let $c_{1}=c_{\min}$ and $c_{K}=c_{\max}.$ As our statistics aggregate over different threshold grids, the problem becomes more demanding. Let $\Delta_{c,\min}=\min_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$ and $\Delta_{c,\max}=\max_{1\leq i\leq{K-1}}|c_{i+1}-c_{i}|$.
and $\underline{d}^{C}=(d_{1K}^{\top},d_{2K}^{\top},\ldots,d_{NK}^{\top})^{\top}.$ Define $Z^{C}=\left({{Z_{1}^{C\top},...,Z_{N}^{C\top}}}\right)^{\top}$ as an $NK$-vector of mean zero Gaussian distributed random vector with covariance matrix $\Sigma^{C}$ specified below. For each $j\in[N]$ and $1\leq i_{1},i_{2}\leq K$, the covariance between the $i_{1}$- and $i_{2}$-th elements of $Z_{j}^{C}$ is given by
Theorem 3.2.
(Gaussian Approximation) Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Suppose that $\Delta_{c,\max}\rightarrow 0.$ Then
$\sup_{u\in\mathbb{R}}\big{|}\mathbb{P}(\mathcal{I}^{C}\leq u)-\mathbb{P}(|Z^{C}+\underline{d}^{C}|_{\infty}\leq u)\big{|}\lesssim\mathcal{R}_{(NK)T},$ (3.7)
The proof of the above theorem is a generalization of Theorem 3.1 with a different variance-covariance structure. It should be noted that the conclusion of the above theorem follows the same pattern as Theorem 3.1, except that $N$ is replaced with $NK$. The following theorem provides the theoretical support for the consistency of the threshold estimator. Note that we only focus on the location $j$ where $\gamma_{j}$ does not equal to 00, and we require those breaks to be significant.
Theorem 3.3 (Consistency of the estimation of thresholds).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Assume $\min_{1\leq j\leq N:\gamma_{j}\neq 0}(Tb_{j})^{1/2}|\gamma_{j}|\gg 1$ and $\Delta_{c,\max}\rightarrow 0.$ Suppose $c_{0j}\in\{c_{i},1\leq i\leq K\}$, then
If the true break $c_{0j}$ does not fall in the grid, then the difference is controlled by $\max\{\Delta_{c,\max},\mathrm{log}(N)/(T\gamma_{j}^{2})\}.$ Under the assumption of a minimum break signal, we can achieve consistency in detecting the break. The rate is expected to depend on both the signal strength $\gamma_{j}$ and the available sample size $T$. Note that the precision of threshold estimation is not affected by the bandwidth $b_{j}$. This rate is in line with the high-dimensional changepoint literature, see for example Theorem 2 in Li et al., (2024).
[…]
4 Monte Carlo Simulations
In this section, we analyze the finite sample performance of our GA results. In particular, we study the size and power of our uniform testing procedures in different settings. The data generating processes(DGPs) 1-7 are presented in Section E in Supplementary Materials.
4.1 Simulation Results
Throughout the study, we use a local linear estimator with a uniform kernel and select the bandwidth according to the procedure described in Calonico et al., (2014). We consider three significance levels, namely, $\alpha=0.1,$ $0.05,$ and $0.01$. The results are based on $1000$ Monte Carlo iterations. All tables and figures are presented in Section E in Supplementary Materials.
The results of simultaneous jump effects testing for DGP1–DGP3 are displayed in Table A1. First, the empirical size closely matches the nominal size across all $N$ values and for large $T$ in all DGPs. For DGP3, we observe slightly oversize when $N=10$ which may be caused by the presence of heteroskedasticity in comparison with results with DGP2. Overall, these simulation results confirm our theoretical findings in Theorem 3.1. In particular, it shows that the test procedure, which is based on critical values neglecting the dependence in the covariance structure, has fairly accurate size control even in the case of strong cross-sectional dependence in the innovations. Second, the local power properties of the test are reasonably well for all DGPs under investigation, and the power increases as either $N$ or $T$ increases. This result confirms that our testing procedure is well-suited to detect sparse alternatives.
In addition, we achieve comparable simulation performance in terms of size and power results for the test of homogeneity of the threshold effects in Table A2. The results confirm our Gaussian approximation result in Theorem A.1. For robustness, we also consider three additional settings with even stronger cross-sectional dependence for DGP4–6 in Table A3. The results show that size and power are indeed robust to strong cross-sectional dependence. Furthermore, Table A4 demonstrates that detecting the existence of jump effects simultaneously across cross-sectional dimensions in the case of unknown thresholds has size and power properties as good as in the known threshold case, confirming our theoretical results from Theorem 3.2. We additionally show the good finite sample performance for the estimation accuracy of the threshold location estimator in Table A5, providing Monte Carlo evidence supporting the consistency result of the threshold in Theorem 3.3.
For comparison, one direct method is the sup-Wald or sup-likelihood ratio test from the pooled threshold regression literature, neither of which accounts for nonlinearity or heterogeneity (see, for example, Andrews, (1993), Hansen, (2000) and Fong et al., (2017)). Notably, recent studies, such as Barassi et al., (2023), propose linear models with heterogeneous threshold effects in panel settings. However, they do not provide uniform testing methods. To facilitate comparison, we firstly restrict ourself to linear threshold model DGP7. And under the alternative we sample a fraction of coefficients $\gamma_{j}$ from a standard normal distribution. As shown in Table A6, our method has similar performance as the sup-Wald type statistic under the null hypothesis, while it has much better power under the alternative. See also Figure 1 in Section 2. The reasons for this are twofold. First, our uniform testing procedure is robust towards settings with sparse signal, in which a pooled test suffers from dilution of signals. And second, our uniform testing procedure does not have the problem of signal cancellation in case the signals for different $j$’s have opposite signs and cancel each other out. As a second comparison, we consider the nonparametric pooled test statistic under DGP3. Again, we draw the non-zero threshold coefficients from a standard normal. Such a setup without accounting for parameter heterogeneity is covered such as in Spokoiny, (1998), Yang and Song, (2014) and Chiou et al., (2018). As shown in Table A7, similar to the linear case, our method works with better power. The above superior performance is mainly due to the sparsity of the signal under the alternative hypothesis. In addition, our method has better power even for dense alternative comparing to the pooled type test statistic when there might exist signal cancellation as illustrated in Figure 1.
We also consider the test of existence of a threshold effect in the derivative under DGP1–3, to achieve similar size and power performance, oversmoothing is needed, as shown in Table A8.
5 Empirical Applications
5.1 Application I: Jumps in the News Impact Curve
In this application, we revisit the news impact curve introduced by Engle and Ng, (1993). In particular, we want to identify possible jumps in the relationship between news events in the form of past returns on the volatility of stocks. The application serves as an illustration of our uniform testing procedure when the threshold $c_{0j}$ is unknown. Originally proposed in the context of autoregressive volatility models (e.g., ARCH), Engle and Ng, (1993) and Linton and Mammen, (2005) consider partially nonparametric specifications for the news impact curve which are able to capture possible asymmetries. A limitation of these specifications is the restriction to continuous functions. However, Zakoian, (1994) and Fornari and Mele, (1997) advocate for (parametric) GARCH specifications that allow for discontinuities in the way past returns impact volatility. We therefore want to investigate this issue by testing for the presence of jumps.
In our model setup, the dependent variable $Y_{jt}$ is the Garman-Klass volatility of stock $j$ at time $t$, and $X_{jt}$ is the lagged stock return. We want to test the null hypothesis $H_{0}^{(1)}:\gamma_{i}=0$ for all $i\in[N]$ and all $c_{0j}$, versus the alternative $H_{a}^{(1)}:\gamma_{i}\neq 0$ for some $i$ and some $c_{0j}$. Our testing procedure is robust to the presence of unobserved risk factors, $U_{jt}$, which are more than likely to occur in our application. Our data includes the daily stock returns and volatility data of the S&P 500 constituents from January 2010 to May 2024, which we obtained from Kaggle  [^2] https://www.kaggle.com/datasets/andrewmvd/sp-500-stocks.. Therefore we have a fairly high-dimensional setting with $N=500$ and $T_{j}$ is the number of days of stock $j$ being a member of the stock index. See Figure 2 for a visualization of the time series of returns and volatility for the stock of AT&T. As candidates for the threshold location, we consider $c_{i}\in\{-0.01,-0.005,0,0.005,0.01\}$.
Figure 2: Time series for daily stock returns (left panel) and daily volatility (right panel) for AT&T. The dashed lines in the left panel indicate the $10\%$ and $90\%$ quantiles of the return distribution.
Figure 3: Local linear fit of a significant threshold effect for AT&T (left panel) and an insignificant effect for Akamai (right panel).
The main findings are summarized in Table A9. We find significant results for $11$ out of the $500$ stocks we consider. For stocks with significant threshold effects, the locations of the estimated thresholds vary, and none of the significant threshold effects occur precisely at zero. This contradicts the rationale of the sign-switching ARCH model of Fornari and Mele, (1997), which models a jump effect based on the sign of the lagged returns, and thus expects a significant threshold effect at zero. For $c_{i}=-0.01$, we get significant negative effects at the $1\%$ level for five stocks, including Allstate Corporation and American Express. On the other hand, we obtain significant positive effects for the $c_{i}=0.01$ threshold for four stocks, e.g. AT&T and McKesson. These results reaffirm the findings of Chen and Ghysels, (2011), who find that both bad and good news have an increasing effect on volatility. Additionally, we get another negatively significant coefficient at the $-0.005$ threshold location and another positively significant coefficient at $c_{i}=0.005$. Overall, we get evidence for the presence of jump effects in the news impact curve. Figure 3 shows a visualization of the significant positive threshold effect for AT&T at $c_{i}=0.01$, and as a comparison we show the fit for Atamai for which we did not find a significant effect at any threshold location.
Furthermore, we are interested in studying the predictive advantage of our method compared to a baseline model without jumps. The baseline model can be written as $Y_{jt}=\tilde{h}_{j}(X_{jt})+e_{jt}$, where $\tilde{h}_{j}(\cdot)$ and $e_{jt}$ are defined in (2.1). For this purpose, we split the data into two parts, with training data from 2010–2019 and test data from 2020–2024. We restrict our analysis to the stocks identified in Table A9 with significant threshold effects at $1\%$ level. The mean squared error of prediction is lower for all but one of the 11 stocks we consider. The same is true if we restrict our analysis to these observations close to the respective threshold, $c_{0j}$. We conclude that including threshold effects can slightly improve the predictive accuracy with a median gain of $1.6\%$ in predictive power.
5.2 Application II: Party Incumbency Effects on U.S. House Elections
In the second application, we are interested in estimating the advantage a candidate has in the U.S. House elections when the seat is currently occupied by the party of the candidate  [^3] We use the data provided by the MIT Election Data and Science Lab, (2017). The data include the vote shares of all candidates in the elections for the US House of Representatives from 1976–2020. Due to redistricting at the start of each decade, we have to exclude years ending with a ‘0’ or ‘2’.. This is called the party incumbency effect. We follow the research design of Lee, (2008). See Caughey and Sekhon, (2011) for an analysis of more recent elections. The dependent variable $Y_{jt}$ is the democratic two-party vote share in year $j$ in district $t$. The covariate $X_{jt}$ is the difference in the two-party vote share in the previous election. Since the winner of the election is determined by the 0 threshold, our model setup is appropriate. We want to test the one-sided null hypothesis $H_{0}^{(1)}:\gamma_{i}<0$ for all $i\in[N]$, versus the alternative hypothesis $H_{a}^{(1)}$: $\gamma_{i}\geq 0$ for some $i$. Unlike Lee, (2008) who pools the data together to estimate a single incumbency effect, we conduct a separate analysis for different election years and also a separate analysis for different states by interchanging the roles of $j$ and $t$ and allowing for possible heterogeneity of incumbency effects along either the $j$ or $t$ dimension. In addition, it is easy to see our theories extend to the unbalanced data in which case we can use $T_{j}$ to denote the number of observations associated with individual/group $j$.
[…]
Now, we are interested in the state-specific effects. We restrict our attention to states with a sufficient number of elections held in the sample. We choose a threshold of 50 elections. This gives us $N=29$ states in our analysis, and $T_{j}$ now denotes the number of elections included for the $j$th state. We argue that this setting constitutes a fairly high-dimensional setting since the effective number of observations is smaller than $N$ for a fairly large proportion of the states. The results are reported in Table A12. Consistent with the previous results, almost all of the estimated coefficients are positive. By using the critical values for the simultaneous tests, we find significant effects for six states, for Missouri at $1\%$ level, and for California, Iowa, Virginia, Alabama and South Carolina at $5\%$. Again, we can reject the null hypothesis of no jump effect for our uniform testing procedure at a $1\%$ level. The significant states tend to belong to those with the largest number of observations. However, for the other two largest states, New York and Texas, the estimated effect is rather small and far from being significant. The incumbency effect seems to be stronger in some states compared to others. See Figure F.4 in the Supplement. for a visualization of the estimated threshold effects for the states of California and Pennsylvania. We also run the test for the homogeneity of incumbency effects ($H_{0}^{\left(2\right)}$) over the 29 states. The results are displayed in Table A13 and the test statistic is given by $\hat{\mathcal{Q}}=3.011$, which implies that we can reject the null hypothesis of homogeneous effects at the $5\%$ level. We find heterogeneity in the effects across election periods and states.
6 Conclusion
This paper focuses on the estimation and inference of heterogeneous threshold effects in nonparametric high-dimensional data, accounting for both cross-sectional and time dependencies in the covariates and error terms. We propose a test to assess the significance of the threshold effect and another to evaluate whether the threshold effects are homogeneous across individuals or groups. Additionally, we introduce a consistent method for estimating unknown threshold locations. Our tests are based on high-dimensional Gaussian approximation results. Despite the complexity of the underlying dependency structures, we show that the variance-covariance structure of the threshold effect estimators has a simple analytical expression under general dependence conditions. Unlike changepoint analysis in time series, our approach considers a setup involving latent variables, leading to a significantly different development of the GA results. Simulations demonstrate significant power enhancement compared to the existing literature under various settings.
SUPPLEMENTARY MATERIALS
Appendix A Additional Main Theorems
A.1 Additional Assumptions
Assumption SA-1.
@@ A.1 Additional Assumptions
Due to the construction of $\varepsilon_{j}(\cdot,\cdot)$ in (2.2), if functions $h_{j}(\cdot,\cdot)$ and $\tau_{j}(\cdot,\cdot)$ are both bounded, then the above assumption naturally holds. The next assumption imposes some smoothness conditions.
As in model (2.1), denote
$f_{j}(x)=\tilde{h}_{j}(x)+\tilde{\tau}_{j}(x)\mathbf{1}_{\{x\geq c_{0j}\}}-\tilde{\tau}_{j}(c_{0j}).$ (SA-1)
Assumption SA-2.
@@
Assumption SA-2 $(i)$ imposes a lower bound condition for the density functions $g_{j}(\cdot)$. Condition $(ii)$ imposes some smoothness conditions of the mean function $f_{j}(\cdot).$ Note that we do not require $f_{j}(\cdot)$ to be differentiable at point $x$; it suffices for $f_{j}(\cdot)$ to have bounded left and right second derivatives near $x$. Assumption $(iii)$ is for the smoothness of the variance function $\sigma_{j}^{2}(\cdot,\cdot)$. If $\sigma_{j}(x,u)$ has bounded partial left and right derivatives with respect to $x$, for any $x\in\mathcal{C}_{j}$ uniformly over $u,$ then assumption $(iii)$ is fulfilled. Note that we do not assume the left and right derivatives of $f_{j}(\cdot)$ to be the same, meaning the slope around the threshold can differ. To be general, we do not impose any differentiability assumptions on $\tilde{h}_{j}(\cdot)$ and $\tilde{\tau}_{j}(\cdot)$.            $\Box$
@@
Assumption SA-3$(i)$-$(ii)$ imposes certain conditions on the kernel function $K\left(\cdot\right)$and bandwidth parameters $b_{j}$’s, respectively. If one chooses $b_{j}$ to be the same across $j,$ Assumption SA-3$(ii)$ reduces to the standard conditions on the bandwidth parameter in local linear kernel estimation with one-dimensional covariate.            $\Box$
A.2 A Homogeneity Test
Theorem A.1 (Gaussian Approximation 2).
Suppose that Assumptions 3.1-3.3 and SA-1-SA-3 hold. Then we have
$\sup_{u\in\mathbb{R}}\Big{|}\mathbb{P}\Big{(}\mathcal{Q}\leq u\Big{)}-\mathbb{P}(|Z+\tilde{\underline{d}}|_{\infty}\leq u)\Big{|}\lesssim\mathcal{R}_{NT}.$ (SA-3)
Theorem A.1 establishes the foundation for testing the null hypothesis $H_{0}^{\left(2\right)}.$ Its proof cannot directly follow from Theorem 3.1 due to the introduction of the additional averaging term of $Y_{jt}$ which might not be of smaller order. However, the steps can be simplified by imposing weak cross-sectional dependence, as the additional averaging term would then be of higher order under this assumption.